Determine the mom Cartesian vector. ent of the for ce Fabout the a-a axis. Express the result as a 2m- 1 m 4 m
Answer #1

Solution: 0 Ma=1-5 20 0 =1-1 50 - 69/5-13.8 kN-m

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solution.